OpenStudy (anonymous):

Surface charge density of a sphere of radius 10 cm is 8.85*10^-8 C/m^2.Find Potential at the centre of the sphere

OpenStudy (experimentx):

hollow sphere or solid sphere??

OpenStudy (anonymous):

Solid

OpenStudy (experimentx):

potential is constant ... i suppose.

OpenStudy (anonymous):

Potential is constant at only the center because E=0 only at the centre

OpenStudy (jamesj):

oh wait ... I'm wrong. E = 0 and hence the potential in the interior must be what it is at the surface.

OpenStudy (anonymous):

But why?E is only 0 at the centre.Jusr near the centre Enot equal to 0

OpenStudy (jamesj):

Yes, it must be because by Gauss's law, the electric flux for any closed surface inside the sphere is zero, because such a surface contains no charge.

OpenStudy (experimentx):

first of all ... let's calculate the charge ... total charge

OpenStudy (experimentx):

are we supposed to assume that charge is only at surface??

OpenStudy (anonymous):

Yes

OpenStudy (jamesj):

that is a fundamental result of electricity, that all charge is at the surface.

OpenStudy (experimentx):

let me open my old book

OpenStudy (anonymous):

All charge is at the surface if and only if it is a conductor.What if it is not

OpenStudy (jamesj):

you're told the sphere has 1 value of electric field density, so even if it actually isn't a conductor, it looks like one right now.

OpenStudy (jamesj):

**electric CHARGE density.

OpenStudy (anonymous):

Well I suppose that may be true

OpenStudy (experimentx):

that would be a great problem in finding the total charge ... but if it were conductor .. our problem will be simplified.

OpenStudy (jamesj):

it's easy: \( Q = \sigma A \)

OpenStudy (anonymous):

Yeah here if it was volume charge density then it would be tough

OpenStudy (jamesj):

and you need this. The electric potential at the surface is \[ V = \frac{kQ}{r} \] and that is the potential in the interior as well.

OpenStudy (anonymous):

Yeah that's right.I was confused because I was thinking what if it was like charge is spread across the volume of the sphere

OpenStudy (experimentx):

yeah i have the same answer

OpenStudy (anonymous):

What is the answer?

OpenStudy (jamesj):

\[ V = \frac{kA\sigma}{r} \] now calculate

OpenStudy (anonymous):

Yeah Im not getting the correct answer

OpenStudy (jamesj):

I guess you could simplify it even one more step to \[ V = 4 \pi k \sigma r \]